ELECTRIC FIELD INTENSITY BETWEEN THE TWO PLATES IS 100N/c ., EACH PLATE OF AREA 2*10 -2 CM 2 . FIND CHARGE ON EACH PLATE...

  • ELECTRIC FIELD INTENSITY BETWEEN THE TWO PLATES IS 100N/c ., EACH PLATE OF AREA  2*10-2  CM2 . FIND CHARGE ON EACH PLATE...
 

Grade:11

1 Answers

Vikas TU
7 years ago
ELECTRIC FIELD INTENSITY BETWEEN THE TWO PLATES = Q/2e0A = 100
Hence,
Q/2Ae0 = 100
Q/Ae0 = 200 
This would be the E in each plate.